a 10m steel rod has a coefficient of linear expansion 1.2 E -2 / Deg C and has is heated from 15 deg C to 45 degree C. Find the change in length of the rod.
Answer #1

Solution) l=lom x = 1.2x102/ec Ti=15°c 72 = 45°C jl=? Al- lLAT la (T2-T ) Al=(10) (1.2 +80°) (us-rs) Al = 3.6